OpenStudy (avountae2003):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept

OpenStudy (cwrw238):

the y intercept can be found by putting x = 0 y = -2(0) - 3 = ? and x intercept by solving for x: -2x - 3 = 0

OpenStudy (cwrw238):

the first one will have coordinates (o , y) and second will be (x , 0)

OpenStudy (avountae2003):

is this right y = -2(0) - 3 = -3 -2x - 3 = 0+ -3/2